# Deep Submillimeter and Radio Observations in the SSA22 Field. IV. Spectral Energy Distributions, Star Formation Histories, and the Infrared-Radio Correlation of the 850 -selected SMGs

### v1

@@ IV.2 Stellar Mass Assembly History
Following the method of Xiao et al. (2024), we place simple constraints on the maximum galaxy mass within the survey volume in a CDM framework. Here, is the baryon-to-star conversion efficiency, is the cosmic baryon fraction, and is the maximum dark matter halo mass. The survey coverage (; Section II) corresponds to a comoving volume of , assuming a conservative redshift range of 0-6. Using the Python package hmf (Murray et al., 2013), we compute the cumulative halo mass function . The maximum detectable halo mass at a given time satisfies . Using the cosmic baryon density and matter density , we adopt a baryon fraction (Aghanim et al., 2020), assumed to be constant. We adopt as the upper limit following common practice (Wechsler and Tinker, 2018), even though that studies suggest the conversion efficiencies range of 20-40% (Behroozi et al., 2019; Girelli et al., 2020; Shuntov et al., 2022).
@@
While the inferred efficiencies do not exceed 100%, such high values remain puzzling. They may stem from overestimated masses or redshifts in SED fitting, or from oversimplified SFH parameterizations, which may vary substantially with environment and mass. Alternatively, it may reflect simplistic assumptions about universal invariants—such as applying a local IMF to high-redshift galaxies, or assuming a constant baryon fraction across halo masses and redshifts (even though deoend on both; Girelli et al., 2020)—among other factors.
@@ IV.2.1 Age and Lifetime of SMGs
To better characterize the typical evolutionary path, we fit the distributions with log-normal functions and analyze the mode, , which represents the most probable value. Here, and denote the mean and standard deviation of the lognormal distribution, with the median of the log-normal distribution given by . The peak of the observed epoch distribution is at 2.63 Gyr (corresponding to ), with fitted parameters and . These massive galaxies began to form extensively around 1.68 Gyr (), with and . The end of the phase peaks at 2.82 Gyr (), with and .
@@ IV.3.1 Contribution to the SFRD
To quantify the contribution of SSA22 deep field SMGs to the cosmic SFRD, we first sum the SFRs of all galaxies within each redshift bin. These values are then corrected for completeness and false detection rates, and divided by the comoving volume corresponding to the survey area and redshift range (i.e., the difference in comoving volumes between the upper and lower redshift boundaries). Given that the deepest 850 observations ( ¡ 1 mJy beam-1) in the central region provide the most complete multi-wavelength coverage, all statistical results discussed in this section are derived from this area. The uncertainties in SFRD estimates are based on Poisson statistics. For consistent comparison with literature values, we apply the following conversions to SFR and stellar mass measurements originally based on different IMFs. Specifically, SFRs and stellar masses derived using a Salpeter (1955) IMF are multiplied by 0.63 and 0.61, respectively, while those based on a Kroupa and Weidner (2003) IMF are scaled by 0.94 and 0.92, respectively (Madau and Dickinson, 2014; Lim et al., 2020b).
@@
From another perspective, our sample at 1 mainly consists of LIRGs, and at 1-2, includes ULIRGs. However, their contribution is only comparable to -selected ULIRGs (Magnelli et al., 2013), and significantly lower than the total SFRD. Several studies on the dust temperature and luminosity of SMGs provide useful insights: 850 -selected SMGs tend to be cooler than -selected sources (Simpson et al., 2017; Dudzevičiūtė et al., 2020; Liao et al., 2024) and exhibit lower average temperatures compared to local and low- samples (Dudzevičiūtė et al., 2020). At fixed luminosity, cooler SMGs show higher submillimeter fluxes (Swinbank et al., 2014; Cunha et al., 2015). This suggests that at 1-2, current surveys miss numerous 850 -faint but hotter galaxies, highlighting the limitations of 850 observations for detecting warm infrared sources at these redshifts and the need for shorter-wavelength surveys.
@@
At 2.5-3.5, the SSA22 SMGs account for 50-60% of the SFRD relative to Madau and Dickinson (2014). For 3-3.5, ALMA HDUF and S2CLS 850 archival data (Bourne et al., 2017; Dunlop et al., 2017) reach a depth of 0.2 mJy (assuming =1.8), revealing that faint submillimeter galaxies contribute significantly to cosmic star formation (Michałowski et al., 2017). Including the overdense region at 3.09 (solid red dot in the figure), our sample contributes 62% of the cosmic SFRD, consistent with these surveys. However, if we exclude the contribution from the overdense regions (open circles, removing sources within the range), SMGs then account for approximately 25% of the SFRD, consistent with the fraction observed at 2–2.5. This indicates that overdense environments or large-scale structures significantly enhance the contribution to the infrared background.
@@
Dunlop et al. (2017); Zavala et al. (2021) further note that at 4, dusty star-forming galaxies become increasingly rare, marking a transition in cosmic star formation activity from a dust-obscured to an unobscured mode. Their contribution to the cosmic SFRD declines sharply, falling to approximately 25–35% at redshifts –6. The SSA22 850 -selected galaxies reflect this trend, with contributions declining rapidly to 20-25% of the cosmic SFRD at 3, consistent with other surveys.
@@
Recently, analyses have been conducted on JWST-selected optically faint galaxies (OFGs), many of which are H-dropouts or similar optical-faint populations (Williams et al., 2024). The majority of OFGs are extended, high-redshift dusty star-forming systems (Wang et al., 2019; Barrufet et al., 2025), with SFRs range from 10 to 100 M⊙ yr-1 and stellar masses that are 1-2 dex lower than those of typical SMGs (Gottumukkala et al., 2024; Williams et al., 2024). OFGs were often entirely missed by HST and lie below the typical detection limit of ALMA (log10(/L⊙) 12) (Williams et al., 2024). At 4-5, OFGs account for only a modest portion of the dust-obscured SFRD ( 26%; brown circles in figure 9). However, their contribution becomes substantia at 5-7, which comparable to previous estimates derived from extrapolated infrared luminosity functions, although “little red dots” (LRDs) dominate the SFRD at 6. The growing role of OFGs implies that past infrared/(sub)millimeter surveys may have missed up to half of the obscured star formation at these redshifts. High-redshift OFGs thus offer a promising avenue to place tighter constraints on cosmic SFRD in the post-reionization era (Williams et al., 2024).
@@ IV.3.2 Contribution to the SMD
On average, SSA22 SMGs contribute approximately 28% to the cosmic SMD (Figure 10). At 4, dust-obscured and unobscured star formation contribute comparably to the cosmic SFRD (Bourne et al., 2017; Dunlop et al., 2017; Zavala et al., 2021). Zavala et al. (2021) suggested that at higher redshifts ( 4-6), the scarcity of dust-rich DSFGs results in a significant deficit of infrared emission relative to UV emission, indicating that unobscured star formation dominates cosmic star formation at these epochs. Despite their low number density at high redshift—approximately 3-410-6 cMpc-3 at 4–6—these rare, massive galaxies contribute 20–25% to the cosmic SFRD and 35–40% to the SMD (Figure 11).
@@
From the early universe to 1–3, the comoving number density of SMGs increases by roughly an order of magnitude, reaching 4-6 10-5 cMpc-3. The SSA22 SMGs reach peak contributions at 2.5–3.5, accounting for 50-60% of the cosmic SFRD (including overdense regions) and a similar fraction of the SMD (Figure 11). Michałowski et al. (2010) reported that during this epoch, SMGs account for more than 60% of the SMD, and at 2-2.5, their survey resolves nearly all of the stellar mass assembled in dusty systems. Long et al. (2023) used numerical models to estimate the contribution of massive DSFGs to the cosmic SMD, finding excellent agreement with observational SMG surveys over the redshift range 2-4.
@@
Zavala et al. (2021) argued that the contribution of DSFGs to the cosmic SFRD remained relatively constant to the present day after reaching its peak (). Moreover, Long et al. (2023) suggest that the contribution of DSFGs to the SMD shows only mild evolution at . However, as shown in Figure 10, our SMG sample—consistent with Michałowski et al. (2010)—exhibits a rapid decline at , with the contribution of SMGs to both the SFRD and SMD dropping sharply to . The contribution from bright SMGs also diminishes quickly after , becoming negligible by this epoch. This trend is further supported by McKinney et al. (2025), who report a sharp drop in the number density of SMGs, indicating that massive dusty galaxies are increasingly difficult to detect in submillimeter surveys at low redshifts. These observations collectively reflect the quenching of star formation in massive SMGs. Dust-obscured star formation—and the associated buildup of stellar mass—has progressively shifted toward lower-mass dusty systems (i.e., downsizing), whose individual contributions to the cosmic SFRD and SMD are small. Additionally, other populations of dusty galaxies not efficiently traced by 850 selection —such as warmer, less luminous, or more compact systems—may dominate obscured star formation at late times but remain underrepresented in current submillimeter surveys. (i.e., wavelength selection effects, section IV.4)
@@ IV.4 Evolution of SMGs
By 4–5, star formation increasingly occurred within dense, dusty environments, surpassing unobscured modes in contribution (Dunlop et al., 2017). This transition is reflected in the rising dust attenuation observed in both UV- and IR-selected galaxies (Paper III), signaling a global shift from an unobscured to a dust-obscured regime (Zavala et al., 2021). Harikane et al. (2022) suggest that galaxy star formation activity may be regulated by the assembly of dark matter halos. During this epoch, a small subset of the most massive galaxies exhibited exceptionally high baryon conversion efficiencies (Xiao et al., 2024; Merlin et al., 2025) (Figure 7).
SMGs undergo intense star formation primarily at 2–5 (Thomas et al., 2005), likely fueled by gas-rich major mergers (Engel et al., 2010; Toft et al., 2014; Chen et al., 2015; Hodge et al., 2025). Nevertheless, morphological and simulation studies indicate that the distribution of galaxy morphologies among SMGs resembles that of the general field population (Swinbank et al., 2010; Casey et al., 2014; McAlpine et al., 2019; Gillman et al., 2024), suggesting their evolutionary pathways may not be fundamentally distinct. Indeed, a significant fraction ( 40–70%) of SMGs appear to be driven by secular processes—such as smooth gas accretion and disk instabilities—rather than violent interactions (Magnelli et al., 2012; Gillman et al., 2024; Chan et al., 2025; Zhang et al., 2025). Observations consistently place SMGs at the high-mass end of the star-forming main sequence (Michałowski et al., 2012; Cunha et al., 2015; Dunlop et al., 2017; Lim et al., 2020b).
@@
In addition to the aforementioned mechanisms of “mass quenching”, “environmental quenching”, and “halo quenching”, AGN feedback also plays a non-negligible role in quenching star formation in massive galaxies. Low-luminosity AGN are commonly detected in massive quiescent galaxies at 2 (Olsen et al., 2013), suggesting active involvement in quenching. Merlin et al. (2025) report strong broad emission lines indicative of AGN activity in 1/3 of their massive quiescent sample. Stellar feedback, too, can act in concert: intense starbursts may heat and expel gas from even deep potential wells on Myr timescales (Merlin et al., 2012). Given that most observed SMGs are dominated by disk structures, morphological quenching (also known as Q-quenching) may occur if the growth of a central bulge stabilizes the disk, thereby suppressing gas fragmentation and star formation (Martig et al., 2009). Furthermore, rapid mass acquisition can also quench galaxy star formation. Oser et al. (2010) showed that in cosmological simulations, stellar mass assembly appears to occur in “two-phase”: early in situ star formation via cold accretion, followed by later structural growth through minor mergers that increase size and mass without reigniting star formation, thereby accelerating quenching.
@@
Finally, we discuss the “downsizing” phenomenon and the scarcity of SMGs detected at 1 in submillimeter surveys. A robust observational trend in galaxy evolution is “downsizing”: massive galaxies complete their star formation early (at high redshift), while lower-mass systems remain actively forming stars to lower redshifts. First articulated by Cowie et al. (1996) to explain low metallicities in high- damped Lyman- systems, this paradigm has been reinforced by numerous studies. Brinchmann and Ellis (2000) showed that most massive galaxies had already quenched by 1, whereas low-mass galaxies continued forming stars. Environmental dependence was highlighted by Thomas et al. (2005), who found earlier quenching in denser regions. Nelan et al. (2005) and Pilyugin and Thuan (2011) observed a “downsizing” trend in metallicity as a function of galaxy mass. Using a galactic hybrid evolutionary tracks model, Firmani and Avila-Reese (2010) successfully reproduced the “downsizing” signature reflected in the stellar mass–halo mass relation at various redshifts. Overall, the downsizing phenomenon appears to be a general pattern in galaxy evolution, manifested in both early- and late-type galaxies, as well as in the red and blue sequences. Why is downsizing so pervasive? It reflects two intertwined processes: (1) rapid, early star formation and quenching in massive systems, and (2) prolonged, lower-level activity in low-mass galaxies that dominates the SFRD at late times. Noeske et al. (2007a) framed this as “staged galaxy formation”: low-mass galaxies start forming stars later and exhibit longer e-folding times for SFR decline, resulting in higher sSFR at low . Massive galaxies, by contrast, experience early, intense bursts—potentially triggered by internal instabilities (Gillman et al., 2024) or enhanced merger rates in dense environments (Thomas et al., 2005)—and assemble their stars rapidly (Neistein et al., 2006), exhausting their gas reservoirs early.
@@
Observationally, SMGs become increasingly elusive at 1. As shown in Section IV.3.1, their contribution to the SFRD drops sharply below 1.5, with LIRGs taking over, and sub-LIRGs dominating at even later times (Magnelli et al., 2013). We attribute this decline to three primary factors: First, quenching and gas exhaustion. In the massive halos () typical of SMGs, baryon conversion efficiency plummets below 2.5, with star formation efficiency (SFE) falling from 10% to just a few percent (Behroozi et al., 2013; Shuntov et al., 2022). Coupled with short gas depletion times ( few hundred Myr), most SMGs quench before 1.5. Moreover, the cosmic molecular gas density peaks at 1.5–2 and declines thereafter, with gas fractions dropping and depletion times increasing rapidly (Liu et al., 2019). Second, the downsizing effect itself: the most extreme, IR-luminous starbursts in massive galaxies fade, leaving lower-mass systems to dominate ongoing star formation (Miller et al., 2015). Third, selection effects due to wavelength and dust temperature. At fixed infrared luminosity, colder dust produces a peak that shifts to longer wavelengths and consequently yields a higher observed 850 flux (Swinbank et al., 2014; Cunha et al., 2015). Thus, 850 surveys preferentially select colder, high-redshift SMGs (Simpson et al., 2017; Dudzevičiūtė et al., 2020; Liao et al., 2024), while missing warmer, low-redshift analogs (see Figure 2 panel a). Indeed, local ULIRGs typically exhibit warmer dust temperatures than high- SMGs (Symeonidis et al., 2013; Clements et al., 2018), and Chen et al. (2021) demonstrated that submillimeter surveys systematically miss galaxies with hotter dust. This is reflected in the statistics: as redshift decreases, the maximum infrared luminosity of galaxies detected at 850 also declines (see Paper III). Consequently, at 1–1.5, current 850 surveys likely miss a substantial population of dust-obscured star formers that are faint in the submillimeter but bright in shorter-wavelength IR bands. This is consistent with the well-established trend that longer-wavelength selections probe higher-redshift populations (Casey et al., 2014; Zavala et al., 2014; Wang et al., 2017), while shorter wavelengths trace lower-redshift obscured activity (Magnelli et al., 2013).

### v2

@@ IV.2 Stellar Mass Assembly History
Following the method of Xiao et al. (2024), we place simple constraints on the maximum galaxy mass within the survey volume in a CDM framework. Here, is the baryon-to-star conversion efficiency, is the cosmic baryon fraction, and is the maximum dark matter halo mass. The survey coverage (; Section II) corresponds to a comoving volume of , assuming a conservative redshift range of 0-6. Using the Python package hmf (murray2013), we compute the cumulative halo mass function . The maximum detectable halo mass at a given time satisfies . Using the cosmic baryon density and matter density , we adopt a baryon fraction (Aghanim et al., 2020), assumed to be constant. We adopt as the upper limit following common practice (Wechsler and Tinker, 2018), even though that studies suggest the conversion efficiencies range of 20-40% (Behroozi et al., 2019; Girelli et al., 2020; Shuntov et al., 2022).
@@
While the inferred efficiencies do not exceed 100%, such high values remain puzzling. They may stem from overestimated masses or redshifts in SED fitting, or from oversimplified SFH parameterizations, which may vary substantially with environment and mass. Alternatively, it may reflect simplistic assumptions about universal invariants—such as applying a local IMF to high-redshift galaxies, or assuming a constant baryon fraction across halo masses and redshifts (even though depend on both; Girelli et al., 2020)—among other factors.
@@ IV.2.1 Age and Lifetime of SMGs
To better characterize the typical evolutionary path, we fit the distributions with log-normal functions and analyze the mode, , which represents the most probable value. Here, and denote the mean and standard deviation of the lognormal distribution, with the median of the lognormal distribution given by . The peak of the observed epoch distribution is at 2.63 Gyr (corresponding to ), with fitted parameters and . These massive galaxies began to form extensively around 1.68 Gyr (), with and . The end of the phase peaks at 2.82 Gyr (), with and .
@@ IV.3.1 Contribution to the SFRD
To quantify the contribution of SSA22 deep field SMGs to the cosmic SFRD, we first sum the SFRs of all galaxies within each redshift bin. These values are then corrected for completeness and false detection rates, and divided by the comoving volume corresponding to the survey area and redshift range (i.e., the difference in comoving volumes between the upper and lower redshift boundaries). Given that the deepest 850 observations ( 1 mJy beam-1) in the central region provide the most complete multi-wavelength coverage, all statistical results discussed in this section are derived from this area. The uncertainties in SFRD estimates are based on Poisson statistics. For consistent comparison with literature values, we apply the following conversions to SFR and stellar mass measurements originally based on different IMFs. Specifically, SFRs and stellar masses derived using a Salpeter (1955) IMF are multiplied by 0.63 and 0.61, respectively, while those based on a Kroupa and Weidner (2003) IMF are scaled by 0.94 and 0.92, respectively (Madau and Dickinson, 2014; Lim et al., 2020b).
@@
From another perspective, our sample at 1 mainly consists of LIRGs, and at 1-2, includes ULIRGs. However, their contribution is only comparable to -selected ULIRGs (Magnelli et al., 2013), and significantly lower than the total SFRD. Several studies on the dust temperature and luminosity of SMGs provide useful insights: 850 -selected SMGs tend to be cooler than -selected sources (Simpson et al., 2017; Dudzevičiūtė et al., 2020; Liao et al., 2024) and exhibit lower average temperatures compared to local and low- samples (Dudzevičiūtė et al., 2020). At fixed luminosity, cooler SMGs show higher submillimeter fluxes (Swinbank et al., 2014; Cunha et al., 2015). This suggests that at lower redshifts ( 1–2), our current survey depth is likely to miss a population of 850 –faint but warmer galaxies with lower dust masses (Section III.2). Detecting such low-dust-mass systems at low redshift requires observations at shorter submillimeter wavelengths; for example, ultra-deep 450 surveys—such as STUDIES (Lim et al., 2020b)—are better suited to uncover this missing population.
@@
At 2.5-3.5, the SSA22 SMGs account for 50-60% of the SFRD relative to Madau and Dickinson (2014). For 3-3.5, ALMA HUDF and S2CLS 850 archival data (Bourne et al., 2017; Dunlop et al., 2017) reach a depth of 0.2 mJy (assuming =1.8), revealing that faint submillimeter galaxies contribute significantly to cosmic star formation. Including the overdense region at 3.09 (solid red dot in the figure), our sample contributes 62% of the cosmic SFRD, consistent with these surveys. However, if we exclude the contribution from the overdense regions (open circles, removing sources within the range), SMGs then account for approximately 25% of the SFRD, consistent with the fraction observed at 2–2.5. This indicates that overdense environments or large-scale structures significantly enhance the contribution to the infrared background.
@@
Dunlop et al. (2017); Zavala et al. (2021); Bouwens et al. (2022) further note that at 4, dusty star-forming galaxies become increasingly rare, marking a transition in cosmic star formation activity from a dust-obscured to an unobscured mode. Their contribution to the cosmic SFRD declines sharply, falling to approximately 25–35% at redshifts –6. The SSA22 850 -selected galaxies reflect this trend, with contributions declining rapidly to 20-25% of the cosmic SFRD at 3, consistent with other surveys.
@@
Optically faint galaxies (OFGs) in ALMA-detected SMG samples typically reside at higher redshifts (Simpson et al., 2014; Dudzevičiūtė et al., 2020) and may be linked to the formation of compact, spheroidal structures in high-redshift massive galaxies (Smail et al., 2021). More recently, leveraging the high sensitivity of JWST, studies have analyzed a population of OFGs. These sources are often undetected by ALMA and include H-band–dropout or similarly extreme optical/near-infrared–faint systems (Williams et al., 2024). The majority of OFGs are extended, high-redshift dusty star-forming systems (Wang et al., 2019; Barrufet et al., 2025), with SFRs range from 10 to 100 M⊙ yr-1 and stellar masses that are 1-2 dex lower than those of typical SMGs (Gottumukkala et al., 2024; Williams et al., 2024). OFGs were often entirely missed by HST and lie below the typical detection limit of ALMA (log10(/L⊙) 12) (Williams et al., 2024). At 4-5, OFGs account for only a modest portion of the dust-obscured SFRD ( 26%; brown circles in figure 9). However, their contribution becomes substantial at 5-7, which comparable to previous estimates derived from extrapolated infrared luminosity functions, although “little red dots” (LRDs) dominate the SFRD at 6. The growing role of OFGs implies that past infrared/(sub)millimeter surveys may have missed up to half of the obscured star formation at these redshifts. High-redshift OFGs thus offer a promising avenue to place tighter constraints on cosmic SFRD in the post-reionization era (Williams et al., 2024).
@@ IV.3.2 Contribution to the SMD
On average, SSA22 SMGs contribute approximately 28% to the cosmic SMD (Figure 10). At 4, dust-obscured and unobscured star formation contribute comparably to the cosmic SFRD (Bourne et al., 2017; Dunlop et al., 2017; Zavala et al., 2021; Bouwens et al., 2022). Zavala et al. (2021) suggested that at higher redshifts ( 4-6), the scarcity of dust-rich DSFGs results in a significant deficit of infrared emission relative to UV emission, indicating that unobscured star formation dominates cosmic star formation at these epochs. Despite their low number density at high redshift—approximately 3-410-6 cMpc-3 at 4–6—these rare, massive galaxies contribute 20–25% to the cosmic SFRD and 35–40% to the SMD (Figure 11).
@@
From the early universe to 1–3, the comoving number density of SMGs increases by roughly an order of magnitude, reaching 4-6 10-5 cMpc-3. The SSA22 SMGs reach peak contributions at 2.5–3.5, accounting for 50-60% of the cosmic SFRD (including overdense regions) and a similar fraction of the SMD (Figure 11). Long et al. (2023) used numerical models to estimate the contribution of massive DSFGs to the cosmic SMD, finding excellent agreement with observational SMG surveys over the redshift range 2-4.
@@
Zavala et al. (2021) argued that the contribution of DSFGs to the cosmic SFRD remained relatively constant to the present day after reaching its peak (). Moreover, Long et al. (2023) suggest that the contribution of DSFGs to the SMD shows only mild evolution at . However, as shown in Figure 10, our SMG sample exhibits a rapid decline at , with the contribution of SMGs to both the SFRD and SMD dropping sharply to . McKinney et al. (2025) report a sharp drop in the number density of SMGs. The contribution from bright SMGs also diminishes quickly after , becoming negligible by this epoch. These observations collectively reflect the quenching of star formation in massive SMGs. Dust-obscured star formation—and the associated buildup of stellar mass—has progressively shifted toward lower-mass dusty systems (i.e., downsizing), whose individual contributions to the cosmic SFRD and SMD are small. Additionally, other populations of dusty galaxies not efficiently traced by 850 selection—such as intrinsically fainter, warmer systems with lower dust masses, or more compact systems—may dominate obscured star formation at late times but remain underrepresented in current submillimeter surveys. (i.e., wavelength selection effects, section IV.4)
@@ IV.4 Evolution of SMGs
By 4–5, star formation increasingly occurred within dense, dusty environments, surpassing unobscured modes in contribution (Dunlop et al., 2017). This transition is reflected in the rising dust attenuation observed in both UV- and IR-selected galaxies (Paper III), signaling a global shift from an unobscured to a dust-obscured regime (Bourne et al., 2017; Dunlop et al., 2017; Bouwens et al., 2020, 2022; Zavala et al., 2021). Harikane et al. (2022) suggest that galaxy star formation activity may be regulated by the assembly of dark matter halos. During this epoch, a small subset of the most massive galaxies exhibited exceptionally high baryon conversion efficiencies (Xiao et al., 2024; Merlin et al., 2025) (Figure 7).
SMGs undergo intense star formation primarily at 2–5 (Thomas et al., 2005), likely fueled by gas-rich major mergers (Tacconi et al., 2008; Engel et al., 2010; Toft et al., 2014; Chen et al., 2015; Hodge et al., 2025). Nevertheless, morphological and simulation studies indicate that the distribution of galaxy morphologies among SMGs resembles that of the general field population (Swinbank et al., 2010; Casey et al., 2014; McAlpine et al., 2019; Gillman et al., 2024), suggesting their evolutionary pathways may not be fundamentally distinct. Indeed, a significant fraction ( 40–70%) of SMGs appear to be driven by secular processes—such as smooth gas accretion and disk instabilities—rather than violent interactions (Magnelli et al., 2012; Gillman et al., 2024; Chan et al., 2025; Zhang et al., 2025). Observations consistently place SMGs at the high-mass end of the star-forming main sequence (Cunha et al., 2015; Dunlop et al., 2017; Lim et al., 2020b).
@@
In addition to the aforementioned mechanisms of “mass quenching”, “environmental quenching”, and “halo quenching”, AGN feedback also plays a non-negligible role in quenching star formation in massive galaxies. Low-luminosity AGN are commonly detected in massive quiescent galaxies at 2 (Olsen et al., 2013), suggesting active involvement in quenching. Merlin et al. (2025) report strong broad emission lines indicative of AGN activity in 1/3 of their massive quiescent sample. Stellar feedback, too, can act in concert: intense starbursts may heat and expel gas from even deep potential wells on Myr timescales (Merlin et al., 2012). Given that most observed SMGs are dominated by disk structures, morphological quenching (also known as Q-quenching) may occur if the growth of a central bulge stabilizes the disk, thereby suppressing gas fragmentation and star formation (Martig et al., 2009). Furthermore, rapid mass acquisition can also quench galaxy star formation. Oser et al. (2010) showed that in cosmological simulations, stellar mass assembly appears to occur in “two-phase”: early star formation via cold accretion, followed by later structural growth through minor mergers that increase size and mass without reigniting star formation, thereby accelerating quenching.
@@
Finally, we discuss the “downsizing” phenomenon and the scarcity of SMGs detected at 1 in submillimeter surveys. A robust observational trend in galaxy evolution is “downsizing”: massive galaxies complete their star formation early (at high redshift), while lower-mass systems remain actively forming stars to lower redshifts. First articulated by Cowie et al. (1996) to explain low metallicities in high- damped Lyman- systems, this paradigm has been reinforced by numerous studies. Brinchmann and Ellis (2000) showed that most massive galaxies had already quenched by 1, whereas low-mass galaxies continued forming stars. Environmental dependence was highlighted by Thomas et al. (2005), who found earlier quenching in denser regions. Nelan et al. (2005) and Pilyugin and Thuan (2011) observed a “downsizing” trend in metallicity as a function of galaxy mass. Using a galactic hybrid evolutionary tracks model, Firmani and Avila-Reese (2010) successfully reproduced the “downsizing” signature reflected in the stellar mass–halo mass relation at various redshifts. Overall, the downsizing phenomenon appears to be a general pattern in galaxy evolution, manifested in both early- and late-type galaxies, as well as in the red and blue sequences. Why is downsizing so pervasive? It reflects two intertwined processes: (1) rapid, early star formation and quenching in massive systems, and (2) prolonged, lower-level activity in low-mass galaxies that dominates the SFRD at late times. Noeske et al. (2007a) framed this as “staged galaxy formation”: low-mass galaxies start forming stars later and exhibit longer e-folding times for SFR decline, resulting in higher sSFR at low redshift. Massive galaxies, by contrast, experience early, intense bursts—potentially triggered by internal instabilities (Gillman et al., 2024) or enhanced merger rates in dense environments (Thomas et al., 2005)—and assemble their stars rapidly (Neistein et al., 2006), exhausting their gas reservoirs early.
@@
Observationally, SMGs become increasingly elusive at 1. As shown in Section IV.3.1, their contribution to the SFRD drops sharply below 1.5, with LIRGs taking over, and sub-LIRGs dominating at even later times (Magnelli et al., 2013). We attribute this decline to three primary factors: First, quenching and gas exhaustion. In the massive halos () typical of SMGs, baryon conversion efficiency plummets below 2.5, with star formation efficiency (SFE) falling from 10% to just a few percent (Behroozi et al., 2013; Shuntov et al., 2022). Coupled with short gas depletion times ( few hundred Myr), most SMGs quench before 1.5. Moreover, the cosmic molecular gas density peaks at 1.5–2 and declines thereafter, with gas fractions dropping and depletion times increasing rapidly (Liu et al., 2019). This is reflected in the statistics: as redshift decreases, the maximum infrared luminosity of galaxies detected at 850 also declines. Second, the downsizing effect itself: the most extreme, IR-luminous starbursts in massive galaxies fade, leaving lower-mass systems to dominate ongoing star formation (Miller et al., 2015). Finally, there is the issue of band selection effects. The selection effect at 850 is tied to dust mass. Consequently, under finite detection limits, 850 surveys may miss sources that have higher dust temperatures and higher infrared luminosities but lower dust masses. Observations also indicate that local or low-redshift ULIRG samples generally exhibit higher dust temperatures than SMGs (Symeonidis et al., 2013; Clements et al., 2018). Moreover, the negative ‑correction at 850 is weaker at lower redshifts; as redshift decreases, the observed flux density at 850 drops rapidly along the Rayleigh–Jeans tail. For a given infrared luminosity, SMGs with cooler dust temperatures show higher flux densities in the submillimeter bands (Swinbank et al., 2014; Cunha et al., 2015). Therefore, 850 observations are more sensitive to the colder dust component (Chapman et al., 2004; Simpson et al., 2017; Dudzevičiūtė et al., 2020; Liao et al., 2024), a bias that becomes particularly pronounced at lower redshifts. This implies that single-band 850 selection may suffer from incompleteness. Within the redshift range 1-1.5, limited by current 850 detection depths, a substantial population of galaxies with relatively high dust temperatures but insufficient dust mass is likely missed. This conclusion is consistent with multiple studies showing that longer-wavelength submillimeter/millimeter observations are more efficient at detecting high-redshift sources (Casey et al., 2014; Zavala et al., 2014; Wang et al., 2017; Casey et al., 2021). Shorter-wavelength far‑infrared/submillimeter bands predominantly probe low-redshift galaxies and contribute more strongly to the low-redshift end of the infrared luminosity function (Magnelli et al., 2013).